# Learning Continually by Spectral Regularization

### v1

@@ Appendix C Additional Experiments
C.1 Other Baselines
Figure 8: AdamW in Random Label MNIST: AdamW performs similarly to L2 regularization towards zero, recovering trainability when LayerNorm is used.
Figure 9: Concatenated ReLU, Random Label MNIST. CReLU is unable to maintain trainability even with layer normalization.
Figure 10: Recycling Dormant Neuron, Random Label MNIST. Recycling dormant neurons are effective without layer normalization, but layer normalization removes dormant neurons and still leads to loss of trainability.
C.2 Ablating Hyperparameter $k$ in Spectral Regularization
Figure 11: Evaluating the choice of $k$ for spectral regularization for $k=1,2,4,8$ (Left) MNIST, (Right) Fashion MNIST.
C.3 Results on Fashion MNIST
Figure 12: Loss of Plasticity in Fashion MNIST Regularization is necessary, LayerNorm without regularization loses plasticity: the performance difference increases as a function of tasks, even with an increasing number of iterations.
C.4 Effect of Regularization on Capacity in a Single Task
@@
C.5 Investigating Effects of Regularization on Neural Network Properties
Figure 14: Singular Values and Weight Norms: Top: SVHN2, Middle: CIFAR10, Bottom: CIFAR100.
Figure 15: Rank and Singular Value Spread: Top: SVHN2, Middle: CIFAR10, Bottom: CIFAR100.
C.6 Investigating Generalization on Continual Imagenet
We investigate the generalization performance of spectral regularization on Continual ImageNet. In Continual ImageNet, each task is to distinguish between two ImageNet classes. We use the network architecture and training protocol used in [17]. We find that all regularization approaches tested achieve high generalization performance across tasks. The methods we compare are spectral regularization (SpectralRegAgent), L2 towards zero (L2Agent), L2 towards initialization (L2InitAgent), recycling dormant neurons (ReDOAgent), and no regularization (BaseAgent).

### v2

@@ Appendix C Additional Experiments
C.1 Ablating Hyperparameter $k$ for Spectral Regularization
Figure 7: Evaluating the choice of $k$ for spectral regularization for $k=1,2,4,8$ on MNIST (Right) and Fashion MNIST (left). We found that $k=2$ balance stability with effectiveness, and use this value throughout our experiments.
C.2 Results on Fashion MNIST
Figure 8: Loss of Plasticity in Fashion MNIST. Although the all the newtorks in this experiment use LayerNorm, loss of plasticity occurs without regularization: the performance decreases as a function of tasks, even with an increasing number of iterations.
C.3 Vision Transformer Cannot Memorize Random Labels
Figure 9: ResNet and Vision Transformer on the random label memorization task using tiny-ImageNet. The Vision Transformer architecture is particularly well-suited to tasks with structure from which generalization is possible. However, we found that its trainability on random labels is lacking. We were unable to get the Vision Transformer to memorize random labels even on a single task in ImageNet.
C.4 Additional Sensitivity Results
Figure 10: Sensitivity analysis on regularization strength Compared to other regularizers, spectral regularization is insensitive to regularization strength while sustaining higher trainability for any given regularization strength.
Robustness to Non-stationarity in Fashion MNIST
Next, we explore whether our proposed regularizers are effective across different types of non-stationarity. To this end, we use EMNIST because the number of classes is large enough that label flipping induces loss of trainability (Elsayed and Mahmood, 2024). In Figure 11, we found that spectral regularization best maintained trainability across different non-stationarities.
Figure 11: Sensitivity to type of non-stationarity on EMNIST. Spectral regularization is able to consistently maintain high trainability across random label assignment, pixel permutations, and label flipping. When applying random label assignment, L2 (init) and L2 (zero) are unable to attain high trainability as compared to spectral regularization.
Varying the Number of Epochs Per Task
Given that the continual learning problem outlined in Section 2 depends on the number of iterations, it may be the case that a higher number of epochs per task can mitigate loss of trainability. In Figure 12, we found that the number of epochs only delays loss of trainability (see Appendix C.2 for results on Fashion MNIST). Even when the number of epochs per task is high enough to reach 100% accuracy on the first task, loss of trainability eventually occurred without regularization. In contrast, when using spectral regularization, loss of trainabililty was consistently mitigated. We additionally found that spectral regularization is particularly effective at maintaining trainability when the number of epochs per tasks is low. The evidence for this is most striking without Layer Norm, in the top row, where spectral regularization was the only method capable of maintaining its trainability. Even though Layer Norm controls the spectral norm through the activations (Kim et al., 2021), adding Layer Norm alone was not enough to mitigate loss of trainability. However, Layer Norm is synergistic with various forms of regularization, with accuracy being improved regardless of the number of epochs per task.
C.5 Effect of Regularization on Capacity in a Single Task
@@
C.6 Investigating Effects of Regularization on Neural Network Properties
Figure 14: Singular values, weight norms, stable ranks and effective gradient diversity: Left: SVHN2, Middle: CIFAR10, Right: CIFAR100.
C.7 Investigating Generalization on Continual Imagenet
We investigate the generalization performance of spectral regularization on Continual ImageNet. In Continual ImageNet, each task is to distinguish between two ImageNet classes. We use the network architecture and training protocol used in Kumar et al. (2023). We find that all regularization approaches tested achieve high generalization performance across tasks. The methods we compare are spectral regularization (SpectralRegAgent), L2 towards zero (L2Agent), L2 towards initialization (L2InitAgent), recycling dormant neurons (ReDOAgent), and no regularization (BaseAgent).
@@ C.7 Investigating Generalization on Continual Imagenet
C.8 Results on Individual DMC Environments
We report mean returns for DMC environments in Figure 16.
Figure 16: Mean and standard error of return for 7 DMC environments.